OpenStudy (anonymous):

Find the average rate of change for the given function from x = 3 to x = 5. −6 −3 3 6

OpenStudy (anonymous):



OpenStudy (anonymous):

i dont know how to do this

OpenStudy (amoodarya):

\[ \frac{\Delta y }{\Delta x }=\frac{ -2-4) }{ 5-3 }\]

OpenStudy (anonymous):

-6/2 = -3